OpenStudy (anonymous):

decker has one allot containing 27% copper and another containing 60% copper. How many pounds of each must he use to make 49 pounds of a third alloy containing 39% copper? Round to the nearest hundredth

OpenStudy (anonymous):

mansonandmax you here for this? because it is not hard

OpenStudy (anonymous):

well it might not be for you but it is for me... if they are that easy for you then IGNORE my posts!!!!!

OpenStudy (anonymous):

that is not the point. i tried replying to the last one but did not get a response. just wanted to make sure you were here

OpenStudy (anonymous):

so we go slow. decker is going to end up with 49 pounds of an alloy that is 39% copper. how much actual copper is there?

OpenStudy (anonymous):

that is, is it clear what 39% of 49 is?

OpenStudy (anonymous):

let x for 27% let y for 60% x+y= 49 y = 49-x 0.27x+0.6y=49*0.39 Substitution y = 49-x 0.27x+0.6( 49-x)=49*0.39 0.27x+29.4-0.6x=19.11 - 0.33x=19.11 - 29.4 x= -10.29/-0.33 x= 33 Substitution x = 33 33+y= 49 y= 49-33 y=16 27% need 33 pounds 60% need 16 pounds if you like my answer,click good answer,TY